OpenStudy (anonymous):

Explain the steps in solving 2.14x - 12.18 = -5.76 and write the solution to the equation

OpenStudy (aem9296):

\[2.14x-12.18 = -5.7\] add 12.18 to both sides \[2.14x = 6.42 \] divide both sides by 2.14 \[\frac{ 2.14x }{ 2.14 } = \frac{ 6.42 }{ 2.14 }\] \[x=3\]

OpenStudy (aem9296):

i ment to say -5.76 but 6.42 is what it equals to when you add 12.18 to both sides